OpenStudy (anonymous):

lim (1/(1-cosx))-(2/x^2)

OpenStudy (atlas):

your question does not have the limit for which you want the solution

OpenStudy (anonymous):

x=0

OpenStudy (atlas):

x->0

OpenStudy (anonymous):

yes

OpenStudy (atlas):

there is a difference i hope u understand that

OpenStudy (anonymous):

its x approaches 0

OpenStudy (anonymous):

first make them one fraction then use l'hospital rule again and again..

OpenStudy (anonymous):

(l'hospital rule 4 times i think)

OpenStudy (atlas):

is the answer infinity......

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the answer is 1/6 but i am trying to get the steps

OpenStudy (anonymous):

yes.. it is 1/6. as i said make it one fraction first then use l'hospital rule for like 4 times

OpenStudy (anonymous):

i tried it and its does not work with me

OpenStudy (atlas):

\[\frac{ 1 }{ 1-cosx } - \frac{ 2 }{ x^2 }\] Is this the question? I am getiing infinity and I am not using L'hospital's rule

OpenStudy (atlas):

\[\frac{ x^2(1+cosx) - 2 \sin^2x }{ x^2\sin^2x }\]

OpenStudy (atlas):

Dividing nr and dr by x^2

OpenStudy (atlas):

oh yeah i got it I am wrong :P

OpenStudy (john_es):

You have to apply L'Hopital rule 4 times, right. Do you know how?